A farmer buys cattle and sheep for an amount of $53800. He buys 100 animals in total. The sheep were $250 each and the cattle were $730 each.How many sheep and how many cattle did he buy?
Answer by jorel1380(3719)

x+y=100
250x+730y=53800
250x+250y=25000
480y=28800
y=60
x=40
The farmer bought 40 sheep and 60 cattle..